OpenStudy (anonymous):

HELP ME PLEASE. find the equation of a parabola whose vertex is on x= -2, latus rectum 8, passing through (-7/8, 4) and whose axis is parallel to the x-axis. the answers are (y-1)^2 = 8(x+2) and (y-7)^2 = 8(x+2). Show it to me step by step. tHANK YOU :)

OpenStudy (anonymous):

NO

OpenStudy (anonymous):

ok fine lets do it

OpenStudy (anonymous):

rectum (god help me) is 8 so we know where the 8 comes from in the answer right?

OpenStudy (anonymous):

we also know it is parallel to the \(x\) axis so the \(y\) is squared

OpenStudy (anonymous):

im so sorry its just that were having an examination today..

OpenStudy (anonymous):

and since the vertex is on \(x=-2\) we got that the \(x\) coordinate of the vertex is \(-2\)

OpenStudy (anonymous):

put it all together and you get \[(y-k)^2=8(x+2)\]

OpenStudy (anonymous):

lol no need to be sorry pretty much like the last one right? put all the facts together, then solve for the number you do not know

OpenStudy (anonymous):

yes yes the last one..

OpenStudy (anonymous):

put \(x=-\frac{7}{8}\) and \(y=4\) and solve for \(k\)

OpenStudy (anonymous):

you are going to get a quadratic, which is why there are two solutions

OpenStudy (anonymous):

\[(y-k)^2=8(x+2)\] \[(4-k)^2=8(-\frac{7}{8}+2)\]

OpenStudy (anonymous):

where m i going to put the points? in the formula? YES. aaah yes. :)

OpenStudy (anonymous):

yes all these that you had pretty much seemed to work the same way

OpenStudy (anonymous):

okay thank you so much!! sorry for disturbing you :) @satellite73

OpenStudy (anonymous):

put together the facts to get as much of the equation as you can, then plug in the numbers to get the ones you do not know

OpenStudy (anonymous):

got it now!! :) thanks again! goodnight!!

OpenStudy (anonymous):

not disturbing me if i didn't want to answer, i could ignore the question good luck on your exam

OpenStudy (anonymous):

Thanks!!

OpenStudy (anonymous):

yw